# XL-DURel: Finetuning Sentence Transformers for Ordinal Word-in-Context Classification

### v1

@@ 1 Introduction
The Ordinal Graded Word-in-Context (OGWiC) task asks to predict the degree of semantic proximity between two word usages on an ordinal scale (Schlechtweg et al., 2025). While it builds on the earlier WiC (Pilehvar and Camacho-Collados, 2019) and GWiC (Armendariz et al., 2020) tasks, it can be distinguished from these by being formulated as an ordinal classification task (Sakai, 2021). This is similar to ranking in that labels are inherently ranked, but also similar to classification in that exact labels have to be predicted for each test instance, instead of merely and ordering of instances as in ranking tasks. State-of-the-art OGWiC models employ pre-trained Language Models like XLM-R (Conneau et al., 2019) to generate contextualized embeddings for the target word in two different contexts and finetune these with loss functions tailored for binary or nominal data such as the contrastive or the cross-entropy loss (e.g. Cassotti et al., 2023; Kuklin and Arefyev, 2025). We conjecture that current state-of-the-art models do not sufficiently exploit the ranking signal provided by OGWiC training data. In this study, we compare several loss functions for classification, regression and ranking tasks to finetune OGWiC models. We manage to outperform previous models on ordinal and binary data with a ranking objective based on angular distance in complex space. By improving performance on the binary and the ordinal formulation of the task through the same model, we pave the way for a unified treatment of WiC modeling. We publish our top-performing model, XL-DURel, which can be employed as highly optimized, fine-grained and multilingual contextualized embedder for word-meaning-related tasks. [^1] https://huggingface.co/sachinn1/xl-durel
@@ 2.1 WiC Task
The challenge of capturing the dynamic semantics of words has led to the development of various evaluation benchmarks. One notable contribution in this area is the WiC task and the corresponding dataset, introduced by Pilehvar and Camacho-Collados (2019). The WiC task is designed to assess context-sensitive word representations by framing it as a binary classification problem. In this task, each instance consists of a target word $w$, and two usages (or sentences) $u_{1}$ and $u_{2}$. The objective is to determine whether the meaning of the target word remains consistent across the two usages. If the meaning is the same, the instance is labeled ‘TRUE’ (or ‘1’) as in pair (1,2):
@@
Performance on this task is usually evaluated with Accuracy. The first WiC dataset by Pilehvar and Camacho-Collados (2019) was constructed from sense-annotated lexical resources such as WordNet (Fellbaum, 2005), VerbNet (Schuler, 2005), and Wiktionary [^2] https://www.wiktionary.org/.
@@ 2.2 GWiC Task
In one of the subtasks, participants were tasked with predicting the absolute similarity rating for each word pair within each context on a scale from 0 to 10. For word pair man and warrior the gold similarity score is 7.88 in (1) and 3.27 in (2). The shared task used the Harmonic Mean of Pearson and Spearman correlations as an evaluation metric. It can thus be interpreted as a mixture of a regression and a ranking task.
@@ 2.3 OGWiC Task
The OGWiC (Ordinal Graded Word-in-Context) task was introduced as part of the CoMeDi shared task (Schlechtweg et al., 2025), focusing on nuanced and interpretable evaluation of word meaning in context. It aims to address the problems of the WiC and GWiC tasks by defining an ordinal classification task requiring participants to exactly reproduce the median annotated label for a word usage pair on a scale from 1 (unrelated) to 4 (closely related). [^3] Find more details on the scale in Appendix A. The pair (1,2) receives label 4 (identical) while pair (1,3) receives the lower label 2 (distantly related):
@@
Two models excelled in the CoMeDi shared task (Choppa, 2024; Kuklin and Arefyev, 2025): XL-LEXEME (Cassotti et al., 2023) builds upon the Sentence Transformer (Reimers and Gurevych, 2019) architecture and employs a bi-encoder framework within a Siamese network. Vectors are initialized with XLM-R (Conneau et al., 2019) and their similarity is directly optimized using a contrastive loss function (Hadsell et al., 2006), which minimizes the distance between embeddings of sentences with the same meaning (label ‘TRUE’) while maximizing the distance between embeddings of different ones (label ‘FALSE’) around a pre-selected margin. At test time, the model predicts the similarity between two usages using the finetuned base model and thresholds it to infer ordinal labels (see Section 4). A similar approach is taken by the DeepMistake model (Arefyev et al., 2021). Vectors are initialized with XLM-R, sentences concatenated and jointly encoded. Then the target word vectors are extracted and jointly fed into a binary classification head. The model is finetuned using the cross-entropy loss. Similar to XL-LEXEME, DeepMistake is trained on binary WiC-like data. At test time, the model predicts the probability of label ‘TRUE’ and thresholds it to infer ordinal labels.
@@
WiC+CoMeDi 299,805 16,667 21,732 as above as above as above
@@ 3.1 Ordinal WiC
We use the OGWiC data provided by the CoMeDi shared task organizers (Schlechtweg et al., 2025), available in starting kit 1.444https://comedinlp.github.io/ The data was sampled from ordinal WiC datasets across multiple languages, including Chinese, English, German, Norwegian, Russian, Spanish and Swedish. (See Table 3 in Appendix B for details.) The data was cleaned in various steps: Initially, instances with fewer than two annotations or those marked with any “Cannot decide” were excluded. Instances with significant annotator disagreement (more than one point on the scale) were also removed. A median judgment was calculated for each instance, retaining only integer medians for task consistency. The data was split by language, with 70% allocated to training, 20% to testing, and 10% to development, ensuring that no target word overlapped between these splits. (See Table 4 in Appendix B for details.)
@@ 3.2 Binary WiC
is a multilingual extension of the original WiC dataset (Pilehvar and Camacho-Collados, 2019), containing over 80,000 sentence pairs across 12 languages: Bulgarian, Chinese, Croatian, Danish, Dutch, Estonian, Farsi, French, German, Italian, Japanese, and Korean. Training data is available for German, French, and Italian, while development and test sets are provided for all 12 languages. Most of the data was automatically extracted from WordNet or Wiktionary sense inventories, without direct human annotation.
@@
comprises approximately 10,000 sentence pairs spanning five languages: Arabic, Chinese, English, French, and Russian. The dataset contains two distinct subtasks: (1) multilingual word-in-context classification within individual languages, and (2) cross-lingual classification involving context sentences from different languages. The dataset is specifically designed to evaluate model performance across both high- and medium-resource language settings. Unlike XL-WiC, which relies on sense inventories, MCL-WiC is entirely human-annotated.
@@
(Adversarial and Multilingual Meaning in Context) contains on average of 13,074 instances per language spanning 14 language pairs and 15 typologically diverse languages, including English, German, Russian, Japanese, Korean, Mandarin Chinese, Arabic, Indonesian, Finnish, Turkish, Basque, Georgian, Urdu, Bengali, and Kazakh. The dataset supports evaluation of word meaning in context both within individual languages and across different languages, with a particular focus on low-resource scenarios. AM2iCo is constructed by automatically extracting cross-lingual word-in-context pairs from Wikipedia, and then filtering them through human validation and adversarial filtering. [^5] Adversarial filtering is a strategy to make a dataset harder and more useful by removing easy examples that models can solve without actually understanding the task.
Cassotti et al. constructed the training set for XL-LEXEME by merging the official training splits from these three datasets. To further augment the training data, they randomly sampled 75% of each dataset’s development data and added it to the training pool. The remaining 25% of the development data was reserved for hyperparameter tuning and validation.
As part of our study, we concatenate the CoMeDi shared task dataset and the XL-LEXEME dataset into a unified resource. For clarity, we refer to the XL-LEXEME dataset concatenation henceforth as “WiC (train/dev)”. We further refer to the CoMeDi shared task data as “CoMeDi (train/dev/test)”. Additionally, we include the original WiC and MCL-WiC test datasets for evaluation in our experiments. We refer to these as “WiC (test)”. Statistics for the final datasets are given in Table 1.
@@ 3.3 Label Mapping
As summarized in Table 1, we apply a systematic label mapping procedure to align the datasets for unified model training and evaluation. Specifically, we transform binary and ordinal labels to cosine-like labels (interval $[0,1]$) if needed for the respective loss function used for training (see Section 4.2). Similarly, we transform ordinal labels to binary labels if needed. As summarized in Table 1, for ordinal-to-cosine mapping, we utilize Min-Max-Scaling to map labels to the interval $[0,1]$. This maps the ordinal labels as follows: 1 $\rightarrow$ 0.0, 2 $\rightarrow\frac{1}{3}$, 3 $\rightarrow\frac{2}{3}$, and 4 $\rightarrow$ 1.0. For binary-to-cosine mapping, we map label 1 (same sense) to cosine label 1.0 to align with annotation level 4 (identical) on the ordinal scale (cf. Table 2 in Appendix A). Binary label 0 (different sense) is mapped to cosine label $\frac{1}{3}$ to align with level 2 (distantly related/polysemy) on the ordinal scale, based on the assumption that most pairs of usages, especially from the same target word, will be semantically related. For ordinal-to-binary mapping, we group ordinal labels 1 and 2 as binary label 0, and labels 3 and 4 as binary label 1, which is motivated by the idea that polysemy is a relation between senses while context variance is a variation within a sense. This mapping is needed in some cases for evaluation. Following the same logic as for binary-to-cosine, for binary-to-ordinal mapping we assign binary label 0 to ordinal label 2, and binary label 1 to ordinal label 4.
[…]
@@ 4.3 Optimization
During training, the parameters of the base model are adjusted in order to minimize the respective loss function from Section 4.2. For all experiments, we keep the following settings constant: we use XLM-RoBERTa-large (Conneau et al., 2019) as our base model and optimize with AdamW. We set the learning rate to $1\times 10^{-5}$, the batch size to $32$, use no weight decay ($0.0$). All other settings are kept at their default values. For Contrastive Loss we set the margin $m=0.5$, and for the ranking-style losses (CoSENT, AnglE Loss) we use the default $\lambda=20$. We train for 10 epochs, with a linear warm-up over 10% of the total training steps. We evaluate at every 25% of an epoch via Average Precision [^7] We also tried Spearman correlation and did not observe a considerable difference in results. (Contrastive Loss) or Spearman correlation (rest) on dev data (see Section 5), and apply early stopping with a patience of 10 evaluations after warm-up and a minimum score improvement threshold of 0.02. The final checkpoint is chosen by highest performance on dev data, where we predict semantic proximity with cosine similarity. [^8] We also experimented with using the angle difference on dev for AnglE Loss, but did not outperform the cosine similarity.
The base model, XLM-RoBERTa-large, contains 561M parameters. All experiments were conducted on a Linux-based server running Fedora 42, equipped with NVIDIA RTX A6000 GPUs (48 GB VRAM per GPU) and dual Intel Xeon CPUs. We utilized a single GPU per run and estimate the computational runtime per model run to be approximately 40–50 GPU hours. [^9] GitHub Copilot was used to assist with coding during the implementation.
@@ 4.4 Thresholding
For all models, we adopt the CoMeDi shared task baseline approach to map cosine similarities to ordinal labels. At test time, similarities are mapped to ordinal labels using three thresholds $\theta$, which are optimized on the dev set by minimizing the following loss function (cf. Choppa, 2024):
@@
  • $\mathbf{y}$ and $\hat{\mathbf{y}}$ are gold labels and predicted cosine similarities respectively,
  • $\alpha$ is Krippendorff’s $\alpha$ and
  • $\hat{\mathbf{y}}_{\theta}$ is a mapping of cosine similarities to discrete labels according to thresholds $\theta$.
Krippendorff’s $\alpha$ is the task evaluation metric (see Section 5). We aim to find optimal values for $\theta$. This threshold optimization is performed on the dev data and separately for each language to account for language-specific distributional differences in cosine similarity scores. It uses the Nelder–Mead simplex method (Nelder and Mead, 1965).
@@ 4.5 Baseline Models
In our experiments, we employ a number of simple baseline models, as described below. All models used thresholding as explained in Section 4.4 for mapping similarities (or probabilities) to ordinal labels.
@@
estimates semantic proximity using the cosine similarity on a non-finetuned SBERT model initialized with XLM-R-Large.
@@
We include the XL-LEXEME result reported in the CoMeDi shared task as a baseline in our evaluation. Notably, it achieved the second-best performance.
@@
Similarly, we include the DeepMistake result reported in the CoMeDi shared task in our evaluation. It achieved the best performance. It uses XLM-R-Large as base model and was finetuned with the cross-entropy loss for binary classification on MCL-WiC train and dev, the Spanish subset of XLWSD (Pasini et al., 2021) and a binarized version of the Spanish DWUG dataset (Zamora-Reina et al., 2022) (see Sections 2.3 and 3). Because Spanish DWUG is part of our test data, we report additional average performance without Spanish in Section 6.

### v2

@@ 1 Introduction
The Ordinal Graded Word-in-Context (OGWiC) task asks to predict the level of semantic proximity between two word usages on an ordinal scale (Schlechtweg et al., 2025). While it builds on the earlier WiC (Pilehvar and Camacho-Collados, 2019) and GWiC (Armendariz et al., 2020) tasks, it can be distinguished from these by being formulated as an ordinal classification task (Sakai, 2021). This is similar to ranking in that labels are inherently ranked, but also similar to classification in that exact labels have to be predicted for each test instance, instead of merely an ordering of instances as in ranking tasks. State-of-the-art OGWiC models employ pre-trained Language Models like XLM-R (Conneau et al., 2019) to generate contextualized embeddings for the target word in two different contexts and finetune these with loss functions tailored for binary or nominal data such as the contrastive or the cross-entropy loss (e.g. Cassotti et al., 2023; Kuklin and Arefyev, 2025). We conjecture that these models do not sufficiently exploit the ranking signal provided by OGWiC training data. In this study, we aim to overcome this limitation by employing loss functions directly optimizing ranking or regression objectives. We compare these against previous models trained with binary classification objectives and manage to outperform the latter on ordinal and binary data with a ranking objective based on angular distance in complex space. By improving performance on the binary and the ordinal formulation of the task through the same model, we pave the way for a unified treatment of WiC modeling. We publish our top-performing model, XL-DURel, which can be employed as highly optimized, fine-grained and multilingual contextualized embedder for word-meaning-related tasks. [^1] https://huggingface.co/sachinn1/xl-durel
@@ 2.1 WiC Task
The challenge of capturing the dynamic semantics of words has led to the development of various evaluation benchmarks. One notable contribution in this area is the Word-in-Context (WiC) task and the corresponding dataset, introduced by Pilehvar and Camacho-Collados (2019). The WiC task is designed to assess context-sensitive word representations by framing it as a binary classification problem. In this task, each instance consists of a target word $w$, and two usages (or sentences) $u_{1}$ and $u_{2}$. The objective is to determine whether the meaning of the target word remains consistent across the two usages. If the meaning is the same, the instance is labeled ‘TRUE’ (or ‘1’) as in pair (1,2):
@@
Performance on this task is usually evaluated with Accuracy. The first WiC dataset by Pilehvar and Camacho-Collados (2019) was constructed from sense-annotated lexical resources such as WordNet (Fellbaum, 2005), VerbNet (Schuler, 2005) and Wiktionary [^2] https://www.wiktionary.org/.
@@ 2.2 GWiC Task
In one of the subtasks, participants were tasked with predicting the absolute similarity rating for each word pair within each context on a scale from 0 to 10. For word pair man and warrior, the gold similarity score is 7.88 in (5) and 3.27 in (6). The shared task used the Harmonic Mean of Pearson and Spearman correlations as an evaluation metric. It can thus be interpreted as a mixture of a regression and a ranking task.
@@ 2.3 OGWiC Task
The Ordinal Graded Word-in-Context (OGWiC) task was introduced as part of the CoMeDi shared task (Schlechtweg et al., 2025), focusing on nuanced and interpretable evaluation of word meaning in context. It aims to address the problems of the WiC and GWiC tasks by defining an ordinal classification task requiring participants to exactly reproduce the median annotated label for a word usage pair on a scale from 1 (unrelated) to 4 (closely related). [^3] Find more details on the scale in Appendix A. For example, the pair (7,8) receives label 4 (identical) while pair (7,9) receives the lower label 2 (distantly related):
@@
Two models excelled in the CoMeDi shared task (Choppa et al., 2025; Kuklin and Arefyev, 2025): XL-LEXEME (Cassotti et al., 2023) builds upon the Sentence Transformers architecture (Reimers and Gurevych, 2019) and employs a bi-encoder framework within a Siamese network. Vectors are initialized with XLM-RoBERTa (XLM-R, Conneau et al., 2019) and their similarity is directly optimized using a contrastive loss function (Hadsell et al., 2006), which minimizes the distance between embeddings of sentences with the same meaning (label ‘TRUE’) while maximizing the distance between embeddings of sentences with different meanings (label ‘FALSE’) around a pre-selected margin. At test time, the model predicts the similarity between two usages using the finetuned base model and thresholds it to infer ordinal labels (see Section 4.4). A similar approach is taken by the DeepMistake model (Arefyev et al., 2021). Vectors are initialized with XLM-R, sentences concatenated and jointly encoded. Then the target word vectors are extracted and jointly fed into a binary classification head. The model is finetuned using the cross-entropy loss. Similar to XL-LEXEME, DeepMistake is trained on binary WiC-like data. At test time, the model predicts the probability of label ‘TRUE’ and thresholds it to infer ordinal labels.
@@
WiC+CoMeDi 299,805 16,668 21,732 as above as above as above
@@ 3.1 Ordinal WiC
We use the OGWiC data provided by the CoMeDi shared task organizers (Schlechtweg et al., 2025), available in starting kit 1.444https://comedinlp.github.io/ The data comprises 71k word usage pairs sampled from ordinal WiC datasets across multiple languages, including Chinese, English, German, Norwegian, Russian, Spanish and Swedish. (See Table 3 in Appendix B for details.) The data was cleaned in various steps: Initially, instances with fewer than two annotations or those marked with any “Cannot decide” were excluded. Instances with significant annotator disagreement (more than one point on the scale) were also removed. A median judgment was calculated for each instance, retaining only integer medians for task consistency. The data was split by language, with 70% allocated to training, 20% to testing, and 10% to development, ensuring that no target word overlapped between these splits. (See Table 4 in Appendix B for details.)
@@ 3.2 Binary WiC
is a multilingual extension of the original WiC dataset (Pilehvar and Camacho-Collados, 2019), containing over 112k sentence pairs across 12 languages: Bulgarian, Chinese, Croatian, Danish, Dutch, Estonian, Farsi, French, German, Italian, Japanese, and Korean. Training data is available for German, French, and Italian while development and test sets are provided for all 12 languages. Most of the data was automatically extracted from WordNet or Wiktionary sense inventories without direct human annotation. The dataset is distributed together with the original English WiC dataset comprising roughly 7K sentence pairs.
@@
(Multilingual and Cross-lingual Word-in-Context Disambiguation) comprises approximately 10k sentence pairs spanning five languages: Arabic, Chinese, English, French, and Russian. The dataset contains data for two distinct subtasks: (i) multilingual WiC classification within individual languages, and (ii) cross-lingual classification comparing sentences from different languages. The dataset is specifically designed to evaluate model performance across both high- and medium-resource language settings. Unlike XL-WiC, which relies on sense inventories, MCL-WiC is entirely human-annotated.
@@
(Adversarial and Multilingual Meaning in Context) contains roughly 196k instances spanning 14 language pairs and 15 typologically diverse languages, including English, German, Russian, Japanese, Korean, Mandarin Chinese, Arabic, Indonesian, Finnish, Turkish, Basque, Georgian, Urdu, Bengali, and Kazakh. The dataset supports evaluation of word meaning in context both within individual languages and across different languages, with a particular focus on low-resource scenarios. AM2iCo is constructed by automatically extracting WiC pairs from Wikipedia, and then filtering them through human validation and adversarial filtering. [^5] Adversarial filtering is a strategy to make a dataset harder and more useful by removing easy examples that models can solve without actually understanding the task.
Cassotti et al. constructed the training set for XL-LEXEME by merging the official training splits from the three above-described datasets. To further augment the training data, they randomly sampled 75% of each dataset’s development data and added it to the training pool. The remaining 25% of the development data was reserved for hyperparameter tuning and validation.
As part of our study, we concatenate the CoMeDi shared task dataset and the XL-LEXEME dataset into a unified resource. For clarity, we refer to the XL-LEXEME dataset concatenation henceforth as “WiC (train/dev)”. We further refer to the CoMeDi shared task data as “CoMeDi (train/dev/test)”. Additionally, we include the original WiC and MCL-WiC test datasets for evaluation in our experiments. We refer to these as “WiC (test)”. Statistics for the final datasets are given in Table 1.666Unintentionally, we skipped 10 files in Cassotti et al.’s training data package, i.e., AM2iCo dev ar-en/bn-en, XL-WiC dev bg-bg/da-da/en-en/et-et/fr-fr/zh-zh, WiC dev en-en, MCL-WiC dev en-en. In total, these are 12,512 instances.
@@ 3.3 Label Mapping
As summarized in Table 1, we apply a systematic label mapping procedure to align the datasets for unified model training and evaluation. Specifically, we transform binary and ordinal labels to cosine-like labels (interval $[0,1]$) if needed for the respective loss function used for training (see Section 4.2). Similarly, we transform ordinal labels to binary labels if needed. As summarized in Table 1, for ordinal-to-cosine mapping, we utilize Min-Max-Scaling to map labels to the interval $[0,1]$. This maps the ordinal labels as follows: 1 $\rightarrow$ 0.0, 2 $\rightarrow\frac{1}{3}$, 3 $\rightarrow\frac{2}{3}$, and 4 $\rightarrow$ 1.0. For binary-to-cosine mapping, we map label 1 (same sense) to cosine label 1.0 to align with annotation level 4 (identical) on the ordinal scale (cf. Table 2 in Appendix A). Binary label 0 (different sense) is mapped to cosine label $\frac{1}{3}$ to align with level 2 (polysemy) on the ordinal scale, based on the assumption that most pairs of usages, especially from the same target word, will be semantically related, e.g. by contiguity or similarity. For ordinal-to-binary mapping, we group ordinal labels 1 and 2 as binary label 0, and labels 3 and 4 as binary label 1, which is motivated by the idea that ordinal label 2 (polysemy) is a relation between senses while ordinal label 3 (context variance) is a variation within a sense (see Appendix A). This mapping is needed in some cases for evaluation. Following the same logic as for binary-to-cosine, for binary-to-ordinal mapping we assign binary label 1 to ordinal label 4, and binary label 0 to ordinal label 2.
BERTBERTpoolingpoolinguuvvcosine-sim(u,v)(u,v)-1 … 1Sentence 1Sentence 2
@@ 4.3 Optimization
During training, the parameters of the base model are adjusted in order to minimize the respective loss function from Section 4.2. For all experiments, we keep the following settings constant: We use XLM-R-large as our base model and optimize with AdamW. We set the learning rate to $1\times 10^{-5}$, the batch size to $32$ and use no weight decay ($0.0$). All other settings are kept at their default values. For Contrastive Loss we set the margin $m=0.5$, and for the ranking losses (CoSENT, AnglE Loss) we use the default $\lambda=20$. We train for 10 epochs, with a linear warm-up over 10% of the total training steps. We evaluate at every 25% of an epoch via Average Precision [^8] We also tried Spearman correlation and did not observe a considerable difference in results. (Contrastive Loss) or Spearman correlation (rest) between cosine similarities and gold labels on dev data (see Section 5). [^9] We also experimented with using the angle difference on dev for AnglE Loss, but did not outperform the cosine similarity. The final checkpoint is chosen by highest performance on dev data.
The base model, XLM-R-large, contains 561M parameters. All experiments are conducted on a Linux-based server running Fedora 42, equipped with NVIDIA RTX A6000 GPUs (48 GB VRAM per GPU) and dual Intel Xeon CPUs. We utilize a single GPU per run and estimate the computational runtime per model run to be approximately 40–50 GPU hours. [^10] GitHub Copilot was used to assist with coding during the implementation.
@@ 4.4 Thresholding
For all models, we adopt the CoMeDi shared task baseline approach to map cosine similarities to ordinal labels. At test time, similarities are mapped to ordinal labels using three thresholds $\theta$, which are optimized on the dev set by minimizing the following loss function (cf. Choppa et al., 2025):
@@
  • $\mathbf{y}$ are gold labels,
  • $\hat{\mathbf{y}}_{\theta}$ are predicted labels according to thresholds $\theta$ on similarity predictions $\hat{\mathbf{y}}$ and
  • $\alpha$ is Krippendorff’s $\alpha$.
Krippendorff’s $\alpha$ is the task evaluation metric (see Section 5). We aim to find optimal values for $\theta$. This threshold optimization is performed on the dev data and separately for each language to account for language-specific distributional differences in similarity scores. It uses the Nelder–Mead simplex method (Nelder and Mead, 1965). (Find induced thresholds in Appendix C.)
@@ 4.5 Baseline Models
In our experiments, we employ a number of simple baseline models, as described below. All models use thresholding as explained in Section 4.4 for mapping similarities to ordinal labels.
@@
uses the cosine similarity on a non-finetuned SBERT model initialized with XLM-R-Large.
@@
is the XL-LEXEME result reported in the CoMeDi shared task. Notably, it achieved the second-best performance. We use this as a baseline in our evaluation.
@@
is the DeepMistake result reported in the CoMeDi shared task. It achieved the best performance. It uses XLM-R-Large as base model and was finetuned with the cross-entropy loss for binary classification on MCL-WiC train and dev, the Spanish subset of XLWSD (Pasini et al., 2021) and a binarized version of the Spanish DWUG dataset (Zamora-Reina et al., 2022) (see Sections 2.3 and 3). Because Spanish DWUG is part of our test data, we report additional average performance without Spanish in Section 6.